Calculate the percent ionization of propionic acid (C2H5COOH) in solutions of each of the following concentrations (Ka is given in Appendix D in the textbook).

-7.96
Answer #1

C2H5COOH(aq) <===> H+(aq) + C2H5COO-(aq)

Ka for propionic acid is 1.3*10^-5

------------------------------------------------------------------------------------

PART-A

1.3 x10^-5 = x^2 /0.0796-x

neglecting x

1.3*10^-5*(0.0796-x ) = X^2

x^2 = 1.3*10^-5*(0.0796)

x^2 =1.0348*10^-6

x= 101072*10^-3

%Ionization = x * 100 / 0.0796

%Ionization =1.0172*10^-3* 100 / 0.0796

%Ionization = 1.277%

------------------------------------------------------------------------------------------

PART-B

1.3 x10^-5 = x^2 /0.0796-x

neglecting x

1.3*10^-5*(0.0205-x ) = X^2

x^2 = 1.3*10^-5*(0.0205)

x^2 =2.665*10^-7

x= 5.1623*10^-4

%Ionization = x * 100 / 0.0205

%Ionization =5.1623*10^-4* 100 / 0.0205

%Ionization = 2.518%

-------------------------------------------------------------------------------